The following two genotypes are crossed: Aa Bb Cc dd Ee x Aa bb Cc Dd Ee. What will be the proportions of the following genotypes among the progeny of this cross? a.AaBbCcddEe

b.AAbbCCddEE
Answer #1

Parents: Aa Bb Cc dd Ee X Aa bb Cc Dd Ee

a. Progeny: Aa Bb Cc dd Ee

From Aa x Aa, probability of Aa is 1/2

From Bb x bb, probability of Bb is 1/2

From Cc x Cc, probability of Cc is 1/2

From dd x Dd, probability of dd is 1/2

From Ee x Ee, probability of Ee is 1/2

Thus proportion of progeny who is AaBbCcddEe is = 1/2 x 1/2 x 1/2 x 1/2 x 1/2 = 1/32

b) Progeny = AAbbCCddEE

From Aa x Aa, probability of AA is 1/4

From Bb x bb, probability of bb is 1/2

From Cc x Cc, probability of CC is 1/4

From dd x Dd, probability of dd is 1/2

From Ee x Ee, probability of EE is 1/4

Thus the proportion of progeny who is AAbbCCddEE is = 1/4 x 1/2 x 1/4 x 1/2 x 1/4 = 1/256